Prev Next

SW&IT Meeting Minutes 2020-08-24

Attendees: Carl, David, Igor, Matthieu, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200810
LR20200810 - Bec, Matthieu 80.0
LR20200810 - Holmberg, Carl 40.0
LR20200810 - Edmeades, David 40.0
LR20200810 - Sola, Igor 40.0
LR20200810 - Hooper, Stephen 80.0
LR20200810 - Bryant, Trey 48.0
LR20200810 - Zhang, Yang 81.0
LR20200810 - Zhang, Xianyu 86.0

News from management

  • Required: UArizona COVID-19 Training, Available Now

Previous Action Items

Owner Action Due Date Completion Date
DE 4D vm / licence? 20200825  

Open topics

  • archive mountpoint woes (ITxxxx)

Team reports

DE

  • Zoom outage
  • Archive mount
  • I haven't heard any updates on from 4D; are there any?
  • VMWare licence renewal
  • JPH printer/scanner
  • DIMM power cycle set up with mountain crew this morning
  • Reconfigure Vault backups to reduce size of VM
  • Windows scripting frustrations re: archive backup, resolved

YZ

  • Previous:
    • Some TMS work (after test data processing/analysis, discussion of new design, wiki work, meeting, etc).
    • Some LBC related discussion.
    • Completed performance review for this year.
    • Safety training mysteriously disappeared?!
  • Next:
    • TMS work (new software development for second test, system spec/operation document).
    • LBC configuration update?
    • AO-Cravat work and meeting.

CH

  • Last week:
    • Supported AGW1 testing, attempting to diagnose out of range errors, brake issues.
      • oacontrol: Added additional logging within src/pmac/setnewpos/setNewPos.c to create 5.2.1. New code wouldn't link to AGW1 or 2 until copying oacontrol.conf from 5.2. Still had issues controling probe. Switched back to oacontrol 5.2.
    • Update LBTI alert script
    • Develop oacontrol error improvements.
  • This week:
    • LBTI alert script: finish tracking down byte vs. str argument error in pyindi.py.
    • Continue AGW1 support, address oacontrol patch issue.

TB

  • Last Week:
    • Changed all plots aside from RMS to small points (got rid of lines - legend OK?)
    • Added Alta series data to composite plots (green line)
    • Some further limited success with Jupyter (can get notebook to behave like command line)
    • Some sensor reading
  • Next:
    • Resolve issue with wind direction w/o needing a loop to iterate through every data point? (workaround?)
    • More sensor reading
    • More work with Jupyter, getting plot to update "inline"?

MB

  • Last week
    • OC, Management
    • PE's written complete, first batch of 1 on 1
    • Misc. meetings, ATT "chair", plan and gathering documentation
    • devel-webui WIP
    • mech interview
  • This week
    • PE's last batch of 1 on 1
    • mech interview #2
    • ATT s/w ctd.
    • DMS websockets, AO, AdSec, etc. integrate demo ui, LBC stiched png preview, max_bytes for ovms
    • Wired registration follow-ups

XZ

  • Last week
    • DX AdSec Fastlink issue
    • SX AdSec restart
    • Meetings
  • This week
    • DX AdSec Flat related issues
    • SX AdSec related
    • Safety training
    • AO support

IS

Last week:

This week:
  • Safety Training
  • LUCI
    • Finish document and configuration for WebIO
  • PIT
    • Finish PIT release.
    • Work on ways to apply an RPM to it, even though the PIT might not be the best example application but it should be very similar to the OT.
Topic revision: r11 - 24 Aug 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback